# Hopf orbits and the first ECH capacity

### v2

@@ Abstract.
We consider dynamically convex (star-shaped) domains in a symplectic vector space of dimension $4$. For such a domain, a “Hopf orbit” is a closed characteristic in the boundary which is unknotted and has self-linking number $-1$. We show that the minimum action among Hopf orbits exists (without any nondegeneracy hypothesis) and defines a symplectic capacity for dynamically convex domains in four dimensions. We further show that this capacity agrees with the first ECH capacity for dynamically convex domains. Combined with a result of Edtmair, this implies that for dynamically convex domains in four dimensions, the first ECH capacity agrees with the cylinder capacity.
@@ 1. Introduction
1.1. Symplectic capacities and the Viterbo conjecture
@@
$\omega_{0}=\sum_{i=1}^{n}dx_{i}\wedge dy_{i}.$
In particular, define the ball
@@
A symplectic capacity defined in dimension $2n$ is a function $c$ from some set of $2n$-dimensional symplectic manifolds (including at least convex domains in $\mathbb{R}^{2n}$) to $[0,\infty]$, such that:
@@
    In particular, if $X$ is a domain in $\mathbb{R}^{2n}$, this means that $c(rX)=r^{2}c(X)$.
@@
The simplest examples of normalized symplectic capacities are the ball capacity $c_{B}$ and the cylinder capacity $c_{Z}$. If $\operatorname{dim}(X)=2n$, then these are defined by
@@
The statements that $c_{B}$ and $c_{Z}$ are normalized symplectic capacities are equivalent to Gromov’s nonsqueezing theorem. In addition, it follows from the definitions that if $c$ is a normalized symplectic capacity defined on $(X,\omega)$, then
@@
The following conjecture, possibly due to Hofer, has apparently been folklore since the 1990s:
Conjecture 1.2 (“strong Viterbo conjecture”).
All normalized symplectic capacities agree on compact convex domains in $\mathbb{R}^{2n}$.
By (1.4), this conjecture is equivalent to the statement that if $X$ is a compact convex domain in $\mathbb{R}^{2n}$, then $c_{B}(X)=c_{Z}(X)$. The conjecture is known for certain special classes of convex domains, for example for $S^{1}$-invariant convex domains by Ostrover, as reviewed in [GHR, Prop. 1.4], for convex domains in a $C^{3}$ neighborhood of a round ball in $\mathbb{R}^{4}$ by Edtmair [gss_edtmair], and more recently for domains in a $C^{2}$ neighborhood of a round ball in $\mathbb{R}^{2n}$ in [1]. In addition, some normalized symplectic capacities are known to agree on convex domains, as we discuss below.
@@
restricts to a contact form $\lambda$ on $\partial X$. The contact form $\lambda$ determines the Reeb vector field $R_{\lambda}$ characterized by $d\lambda(R_{\lambda},\cdot)=0$ and $\lambda(R_{\lambda})=1$. A Reeb orbit is a periodic orbit of $R_{\lambda}$, i.e. a map $\gamma:\mathbb{R}/T\mathbb{Z}\to\partial X$ for some $T>0$, modulo precomposition with translations of $\mathbb{R}/T\mathbb{Z}$, such that $\gamma^{\prime}(t)=R_{\lambda}(\gamma(t))$ for all $t$. The Reeb orbit $\gamma$ is simple if the map $\gamma$ is an embedding. The symplectic action $\mathcal{A}(\gamma)>0$ is defined to be the period $T$. By a theorem of Rabinowitz [34], at least one Reeb orbit exists. Let $\mathcal{P}(\partial X)$ denote the set of Reeb orbits in $\partial X$, and define
@@
The Ekeland-Hofer-Zehnder capacity is a normalized symplectic capacity $c_{\operatorname{EHZ}}$ which is defined for convex domains in $\mathbb{R}^{2n}$ as follows. If $X$ is also a star-shaped domain as above, then
(1.5) $c_{\operatorname{EHZ}}(X)=\mathcal{A}_{\operatorname{min}}(X).$
It is a nontrivial fact that $\mathcal{A}_{\operatorname{min}}$ is monotone under symplectic embeddings, and then it extends by $C^{0}$ continuity to all convex domains [^2] The history is that the previously defined Ekeland-Hofer and Hofer-Zehnder capacities were shown to agree with the minimal action $\mathcal{A}_{\operatorname{min}}$ for convex star-shaped domains; see [GHR, Thm. 1.12]..
@@
If $X\subset\mathbb{R}^{2n}$ is a star-shaped domain, then for each $k$ there is a finite set of Reeb orbits $\{\gamma_{i}\}$ on $\partial X$, selected homologically by embedded contact homology, such that
$c_{k}^{\operatorname{ECH}}(X)=\sum_{i}\mathcal{A}(\gamma_{i}).$
The first ECH capacity $c_{1}^{\operatorname{ECH}}$ is normalized, and we will review its definition for star-shaped domains in §3.2.
There are also “alternative ECH capacities” $c_{k}^{\operatorname{Alt}}$ defined by a more elementary max-min formula in [28], which satisfy the same basic properties as those of the ECH capacities listed above and agree with them on basic examples. If $X$ is a star-shaped domain, then by [28, Thm. 12] we have
@@
Some additional sequences of symplectic capacities, defined for star-shaped domains in $\mathbb{R}^{2n}$, are the Ekeland-Hofer capacities $\{c_{k}^{\operatorname{EH}}\}_{k\geq 1}$ defined in [11] using variational methods, and the conjecturally equal capacities $\{c_{k}^{\operatorname{CH}}\}_{k\geq 1}$ defined in [13] using equivariant symplectic homology. The capacities $c_{1}^{\operatorname{EH}}$ and $c_{1}^{\operatorname{CH}}$ are normalized. As reviewed in [GHR, Thm. 1.12], if $X$ is a convex domain in $\mathbb{R}^{2n}$ then it is known that
$c_{\operatorname{EHZ}}(X)=c_{1}^{\operatorname{EH}}(X)=c_{1}^{\operatorname{CH}}(X).$
For a convex domain in $\mathbb{R}^{4}$ one can further show that the above capacities agree with $c_{1}^{\operatorname{Alt}}(X)$. However it is not known whether the latter agrees with $c_{1}^{\operatorname{ECH}}(X)$; see [28, Rem. 13].
Remark 1.4.
The original Viterbo conjecture from [viterbo] asserts that if $c$ is a normalized symplectic capacity, and if $X$ is a compact convex domain in $\mathbb{R}^{2n}$, then
(1.6) $c(X)^{n}\leq n!\operatorname{vol}(X),$
with equality if and only if $\operatorname{int}(X)$ is symplectomorphic to an open round ball. Note that the inequality (1.6) follows from the strong Viterbo conjecture. This is because if there exists a symplectic embedding $B^{2n}(r)\hookrightarrow X$, then
$\frac{r^{n}}{n!}=\operatorname{vol}(B^{2n}(r))\leq\operatorname{vol}(X),$
and it follows that $c_{B}(X)^{n}\leq n!\operatorname{vol}(X)$.
Remark 1.5.
The “weak Viterbo conjecture” asserts that if $X$ is a convex star-shaped domain, then $\mathcal{A}_{\operatorname{min}}(X)^{n}\leq n!\operatorname{vol}(X)$. This follows from the inequality in the Viterbo conjecture (1.6) and equation (1.5) since $c_{\operatorname{EHZ}}$ is a normalized symplectic capacity. The weak Viterbo conjecture is known to imply the Mahler conjecture in convex geometry [3]. By continuity, the weak Viterbo conjecture is equivalent to the statement that if $X$ is a convex polytope, then $c_{\operatorname{EHZ}}(X)^{n}\leq n!\operatorname{vol}(X)$. The latter conjecture has been tested numerically in [6, 15].
@@ 1.2. The Hopf capacity
Let $X\subset\mathbb{R}^{4}$ be a star-shaped domain and let $\lambda$ denote the contact form on $\partial X$. The contact structure $\xi=\operatorname{Ker}(\lambda)$ has, up to homotopy, a unique symplectic trivialization $\tau$ over $\partial X$. If $\gamma$ is a Reeb orbit, then the linearized Reeb flow along $\gamma$ has a well-defined rotation number with respect to $\tau$, which we denote by $\rho(\gamma)\in\mathbb{R}$; see [27, §1.2] for details.
@@
The notion of dynamical convexity was introduced by Hofer-Wysocki-Zehnder [convex], as a symplectomorphism-invariant alternative to the notion of convexity. It is shown in [convex] that every strictly convex star-shaped domain in $\mathbb{R}^{4}$ is dynamically convex. For a long time it was not known whether, conversely, every dynamically convex domain is symplectomorphic to a convex domain. It was recently shown by Chaidez-Edtmair [5] that the answer is no; see [9, 8] for further developments. In addition, it is shown in [ABHS_comp] that the conclusion of the weak Viterbo conjecture fails for some dynamically convex domains in $\mathbb{R}^{4}$. Nonetheless, the results that we will prove in this paper are valid for dynamically convex domains.
@@
A Hopf orbit is a simple Reeb orbit $\gamma$ in $\partial X$ which is unknotted and has self-linking number $-1$. Let $\mathcal{P}_{\operatorname{Hopf}}(\partial X)$ denote the set of Hopf orbits in $\partial X$.
@@
It is shown in [convex] that if $X$ is dynamically convex, then the Reeb flow on $\partial X$ admits a disk-like global surface of section. This is an embedded disk $D\subset\partial X$ such that $\partial D$ is a simple Reeb orbit $\gamma$; the Reeb flow is transverse to $\mathrm{int}(D)$; and for every $x\in\partial X\setminus\gamma$, the Reeb trajectory through $x$ intersects $\mathrm{int}(D)$ forwards and backwards in time. It follows from these conditions that $\gamma$ is a Hopf orbit. Conversely, it is shown in [hryn_jsg] that if $X$ is dynamically convex, then every Hopf orbit in $\partial X$ bounds a disk-like global surface of section for the Reeb flow.
@@
$\mathcal{A}_{\rm Hopf}(X)=\inf_{\gamma\in\mathcal{P}_{\rm Hopf}(\partial X)}\mathcal{A}(\gamma).$
@@
(1.8) $c_{\rm Hopf}(X,\omega)\ =\ \sup\{\mathcal{A}_{\rm Hopf}(\Omega)\mid\Omega\text{ is dynamically convex and }\Omega\hookrightarrow(X,\omega)\}$
where the supremum is taken over all dynamically convex domains $\Omega\subset\mathbb{R}^{4}$ such that $\Omega$ symplectically embeds into $(X,\omega)$.
@@
for some $\gamma\in\mathcal{P}_{\rm Hopf}(X)$ satisfying ${\rm CZ}(\gamma)=3$.
Let $X\subset\mathbb{R}^{4}$ be a dynamically convex domain. It follows from Theorem 1.12 and (1.4) that
$\mathcal{A}_{\rm Hopf}(X)\leq c_{Z}(X).$
On the other hand, Edtmair [gss_edtmair] showed that if there exists a Hopf orbit $\gamma$ with $\mathcal{A}(\gamma)=R$, then there exists a symplectic embedding $X\hookrightarrow Z(R)$. In particular, $c_{Z}(X)\leq\mathcal{A}_{\operatorname{Hopf}}(X)$. Combining this with the second part of Theorem 1.12, we get:
Corollary 1.13.
If $X\subset\mathbb{R}^{4}$ is a dynamically convex domain, then
(1.9) $c_{Z}(X)=\mathcal{A}_{\rm Hopf}(X),$
and the infimum in (1.3) is realized.
@@ 1.3. The first ECH capacity
One of the consequences of Theorems 1.12 and 1.14 and their proofs is that we can prove that $c_{1}^{\rm ECH}$ is a symplectic capacity for dynamically convex domains, without using Seiberg-Witten theory as in the proof of monotonicity of ECH capacities in [hut_qech].
@@
To summarize the situation with the strong Viterbo conjecture, for a convex domain $X\subset\mathbb{R}^{4}$, we now know that the normalized symplectic capacities discussed above can be sorted in nondecreasing order as follows:
  • The ball capacity $c_{B}$.
  • The capacities $c_{\operatorname{EHZ}}=c_{1}^{\operatorname{EH}}=c_{1}^{\operatorname{CH}}=c_{1}^{\operatorname{Alt}}$.
  • The capacities $c_{\operatorname{Hopf}}=c_{1}^{\operatorname{ECH}}=c_{Z}$. (These equalities hold more generally in the dynamically convex case.)
Acknowledgments. U.H. acknowledges the support by DFG SFB/TRR 191 ‘Symplectic Structures in Geometry, Algebra and Dynamics’, Projektnummer 281071066-TRR 191. M.H. was partially supported by NSF grant DMS-2005437. V.G.B.R. is grateful for the hospitality of the Institute for Advanced Study, where part of this work was completed. V.G.B.R. was partially supported by NSF grant DMS-1926686, FAPERJ grant JCNE E-26/201.399/2021 and a Serrapilheira Institute grant.

### v3

@@ Abstract.
We consider dynamically convex star-shaped domains in a symplectic vector space of dimension $4$. For such a domain, a “Hopf orbit” is a closed characteristic in the boundary which is unknotted and has self-linking number 1-1. We show that the minimum action among Hopf orbits exists and defines a symplectic capacity for dynamically convex star-shaped domains. We further show that this capacity agrees with the first ECH capacity for such domains. Combined with a result of Edtmair [13], this implies that for dynamically convex star-shaped domains in four dimensions, the first ECH capacity agrees with the cylinder capacity. This also provides a method to show that the first ECH capacity of a dynamically convex star-shaped domain satisfies the axioms of a normalized symplectic capacity without any need for Seiberg-Witten theory.
@@ 1. Introduction
The theme of this paper is the study of symplectic capacities of dynamically convex star-shaped domains in a $4$-dimensional symplectic vector space. In order to state our results and related questions we first need to recall a few basic notions.
1.1. Symplectic capacities
@@
$\omega_{0}=\sum_{j=1}^{n}dx_{j}\wedge dy_{j}.$
Here $(x_{1},\dots,x_{n},y_{1},\dots,y_{n})$ denote the coordinates in $\mathbb{R}^{2n}$. The corresponding coordinates in $\mathbb{C}^{n}$ are denoted by $(z_{1},\dots,z_{n})$, where $z_{j}=x_{j}+iy_{j}\in\mathbb{C}$. In particular, define the ball
@@
A symplectic capacity defined in dimension $2n$ is a function $c$ from some nonempty set of $2n$-dimensional symplectic manifolds to $0,\infty]$, such that:
@@
    In particular, if $X$ is a domain in $\mathbb{R}^{2n}$ then $c(rX)=r^{2}c(X)$ for all $r\in\mathbb{R}\setminus\{0\}$.
@@
The simplest examples of normalized symplectic capacities are the Gromov width $c_{B}$, sometimes also referred to as the ball capacity, and the cylinder capacity $c_{Z}$. If $\operatorname{dim}(X)=2n$, then these are defined by
@@
The statements that $c_{B}$ and $c_{Z}$ are normalized symplectic capacities are equivalent to Gromov’s nonsqueezing theorem [gromov]. In addition, it follows from the definitions that if $c$ is a normalized symplectic capacity defined on $(X,\omega)$, then
@@
restricts to a contact form $\lambda$ on $\partial X$. The contact form $\lambda$ determines the Reeb vector field $R_{\lambda}$ characterized by $d\lambda(R_{\lambda},\cdot)=0$ and $\lambda(R_{\lambda})=1$. A Reeb orbit is a periodic orbit of $R_{\lambda}$, i.e. a map $\gamma:\mathbb{R}/T\mathbb{Z}\to\partial X$ for some $T>0$, modulo translations of $\mathbb{R}/T\mathbb{Z}$, such that $\dot{\gamma}(t)=R_{\lambda}(\gamma(t))$ for all $t$. The Reeb orbit $\gamma$ is simple if the map $\gamma$ is an embedding, i.e. $T>0$ is the smallest positive period. The symplectic action $\mathcal{A}(\gamma)>0$ is defined to be the period $T=\int_{\gamma}\lambda$. By a theorem of Rabinowitz [40], at least one Reeb orbit exists. Let $\mathcal{P}(\partial X)$ denote the set of Reeb orbits in $\partial X$, and define
@@
The action spectrum of the star-shaped domain $X$ is defined to be the set of periods of Reeb orbits in $\partial X$. A symplectic capacity $c$ is said to be spectral on star-shaped domains if for every star-shaped domain $X$ we have
(1.5) $c(X)=\sum_{i}\mathcal{A}(\gamma_{i})$
for some finite collection $\{\gamma_{i}\}$ of Reeb orbits in $\partial X$.
@@
For each $k$ the capacity $c_{k}^{\rm ECH}$ is spectral on star-shaped domains, i.e. if $X\subset\mathbb{R}^{4}$ is a star-shaped domain then (1.5) holds, where the finite set of Reeb orbits $\{\gamma_{i}\}$ on $\partial X$ is selected homologically by embedded contact homology. The first ECH capacity $c_{1}^{\operatorname{ECH}}$ is normalized, and we will review its definition for star-shaped domains in §3.2.
There are also “alternative ECH capacities” $c_{k}^{\operatorname{Alt}}$ defined by a more elementary max-min formula in [34], which satisfy the same basic properties as those of the ECH capacities listed above and agree with them on basic examples. If $X$ is a star-shaped domain, then by [34]*Thm. 12 we have
@@
Some additional sequences of spectral symplectic capacities, defined for star-shaped domains in $\mathbb{R}^{2n}$, are the Ekeland-Hofer capacities $\{c_{k}^{\operatorname{EH}}\}_{k\geq 1}$ defined in [14] using variational methods. Recent work in [18] establishes that these capacities are equal to the capacities $\{c_{k}^{\operatorname{CH}}\}_{k\geq 1}$ defined in [16] using equivariant symplectic homology. The capacity $c_{1}^{\operatorname{EH}}=c_{1}^{\operatorname{CH}}$ is normalized.
@@ 1.2. The Hopf capacity
Let $X\subset\mathbb{R}^{4}$ be a star-shaped domain and let $\lambda$ denote the contact form on $\partial X$. The contact structure $\xi=\operatorname{Ker}(\lambda)$ has, up to homotopy, a unique symplectic trivialization $\tau$ over $\partial X$. If $\gamma$ is a Reeb orbit, then the linearized Reeb flow along $\gamma$ has a well-defined rotation number with respect to $\tau$, which we denote by $\rho(\gamma)\in\mathbb{R}$; see [33, §1.2] for details.
@@
The notion of dynamical convexity was introduced by Hofer, Wysocki and Zehnder in [convex], as a symplectomorphism-invariant alternative to the notion of convexity. It is shown in [convex] that every convex star-shaped domain in $\mathbb{R}^{4}$ whose boundary is positively curved is dynamically convex. For a long time it was not known whether, conversely, every dynamically convex domain is symplectomorphic to a convex domain. It was recently shown by Chaidez and Edtmair in [8] that the answer is no; see [12, 11] for further developments.
@@
A Hopf orbit is a simple Reeb orbit $\gamma$ in $\partial X$ which is unknotted and has self-linking number 1-1. Let $\mathcal{P}_{\operatorname{Hopf}}(\partial X)$ denote the set of Hopf orbits in $\partial X$.
@@
It is shown in [convex] that if $X$ is dynamically convex, then the Reeb flow on $\partial X$ admits a disk-like global surface of section. This is an embedded disk $D\subset\partial X$ such that $\partial D$ is a Reeb orbit $\gamma$, the Reeb flow is transverse to $\mathrm{int}(D)$, and for every $x\in\partial X\setminus\gamma$, the Reeb trajectory through $x$ intersects $\mathrm{int}(D)$ forwards and backwards in time. It follows from these conditions that $\gamma$ is a Hopf orbit. Conversely, it is shown in [hryn_jsg] that if $X$ is dynamically convex, then every Hopf orbit in $\partial X$ bounds a disk-like global surface of section for the Reeb flow.
@@
$\mathcal{A}_{\rm Hopf}(X)=\inf\ \{\mathcal{A}(\gamma)\mid\gamma\in\mathcal{P}_{\rm Hopf}(\partial X)\}.$
@@
(1.7) $c_{\rm Hopf}(X,\omega)=\sup\{\mathcal{A}_{\rm Hopf}(\Omega)\mid\Omega\text{ is dynamically convex, }(\Omega,\omega_{0})\hookrightarrow(X,\omega)\}$
where the supremum is taken over all dynamically convex star-shaped domains $\Omega\subset\mathbb{R}^{4}$ such that $(\Omega,\omega_{0})$ symplectically embeds into $(X,\omega)$.
@@
for some $\gamma\in\mathcal{P}_{\rm Hopf}(X)$ satisfying $\operatorname{CZ}(\gamma)=3$.
Remark 1.10.
Theorem 1.9 and Remark 1.6 relate to the following important question posed by Hofer, Wysocki and Zehnder in [convex]: For a dynamically convex star-shaped domain $X$ in $\mathbb{R}^{4}$, is $\mathcal{A}_{\rm min}(X)$ equal to the action of a Reeb orbit that bounds a disk-like global surface of section for the Reeb flow on $\partial X$? In view of the result from [hryn_jsg], this is equivalent to asking: Is $\mathcal{A}_{\rm min}(X)=\mathcal{A}_{\rm Hopf}(X)$ for every dynamically convex star-shaped domain $X$ in $\mathbb{R}^{4}$? This difficult question remains open. Abbondandolo, Edtmair and Kang showed in [5] that it has an affirmative answer in the special case where $X$ is convex.
Remark 1.11.
The key technical step in the proof of Theorem 1.9 is Lemma 2.1 below. Its proof is based on the notion of a “fast plane” introduced in [27]. An end of a finite-energy curve can be seen as a gradient trajectory of the action functional on the loop space; it converges to a periodic orbit, i.e. a critical point of the action. There is an associated asymptotic operator, which plays the role of the “Hessian” of the action at the critical point; in a nondegenerate situation it governs the approach to the periodic orbit, in the very same way that in finite-dimensions the Hessian governs how a gradient trajectory approaches a critical point. This is the main result of the work [props1] by Hofer, Wysocki and Zehnder. Still assuming nondegeneracy, a curve approaches the periodic orbit at a given puncture with an exponential decay dictated by a negative eigenvalue. The more negative the eigenvalue, the faster the approach. In $3$-dimensions, eigenvalues can be compared to each other in terms of winding numbers of associated eigenvectors: smaller winding number implies smaller eigenvalue. Moreover, for topological reasons, the winding number of the “asymptotic eigenvector” of a finite-energy plane is $\geq 1$ in a trivialization of the contact structure that extends over the plane. A fast plane is one for which the approach is as fast as possible, in the sense that the corresponding eigenvalue has eigenvectors of winding number $1$. It turns out that fast planes have good intersection-theoretic properties to form foliations transverse to the Reeb flow and, under the assumption of dynamical convexity, to induce global surfaces of section.
@@ 1.3. The first ECH capacity
It is shown in [hut_qech] using Seiberg-Witten theory that $c_{1}^{\rm ECH}$ is a normalized symplectic capacity. It follows from this and Theorem 1.12 that $c_{\rm Hopf}$ is a normalized symplectic capacity on dynamically convex star-shaped domains in $\mathbb{R}^{4}$. However our proofs of Theorems 1.9 and Theorem 1.12 establish that $c_{1}^{\rm ECH}$ and $c_{\rm Hopf}$ are normalized symplectic capacities on dynamically convex star-shaped domains in $\mathbb{R}^{4}$ without using Seiberg-Witten theory.
@@
1.4. Relation to the cylinder capacity
Let $X\subset\mathbb{R}^{4}$ be a dynamically convex domain. It follows from Theorem 1.9 or Theorem 1.12 and (1.4) that
(1.8) $\mathcal{A}_{\rm Hopf}(X)\leq c_{Z}(X).$
On the other hand, Edtmair [13] showed that if there exists a Hopf orbit $\gamma$ with $\mathcal{A}(\gamma)=R$, then there exists a symplectic embedding $X\hookrightarrow Z(R)$. In particular, $c_{Z}(X)\leq\mathcal{A}_{\operatorname{Hopf}}(X)$. Combining this with Theorem 1.9, or alternatively with Theorem 1.12, we get:
Corollary 1.15.
If $X\subset\mathbb{R}^{4}$ is a dynamically convex domain, then
(1.9) $c_{Z}(X)=\mathcal{A}_{\rm Hopf}(X),$
and the infimum in (1.3) is realized.
Remark 1.16.
Corollary 1.15 is also stated in [13]. However, it should be stressed that the work done in [13] concerning (1.9) consists solely in the proof of the inequality $c_{Z}(X)\leq\mathcal{A}_{\operatorname{Hopf}}(X)$ for a dynamically convex star-shaped domain $X$. The reverse inequality $\mathcal{A}_{\operatorname{Hopf}}(X)\leq c_{Z}(X)$ is a direct consequence of Theorem 1.9. This is used in [13].
1.5. Viterbo’s conjecture and the convex case
Viterbo’s conjecture from [viterbo] asserts that if $c$ is a normalized symplectic capacity and $X\subset\mathbb{R}^{2n}$ is convex then $c(X)^{n}\leq n!\,{\rm vol}(X)$. This would follow from a positive answer to a question due to Hofer from the 1990s: Do the values of all normalized symplectic capacities agree on a convex domain? A conjectural positive answer to this question has been later called the “strong Viterbo conjecture”. Its validity implies the “weak Viterbo conjecture” asserting that $\mathcal{A}_{\rm min}(X)^{n}\leq n!\,{\rm vol}(X)$ when $X$ is a convex domain with smooth boundary. See [GHR] for discussion. Recently Haim-Kislev and Ostrover  [6] found counterexamples to the weak Viterbo conjecture, and hence to all versions of the conjecture, for all $n\geq 2$. On the other hand, the result of [5] shows that if $X\subset\mathbb{R}^{4}$ is convex, then all normalized capacities of $X$ that are spectral agree with $\mathcal{A}_{\rm min}(X)$. It is an open question whether this is still true if $X\subset\mathbb{R}^{4}$ is dynamically convex. For example, if $X\subset\mathbb{R}^{4}$ is dynamically convex, then we know that $c_{1}^{\rm Alt}(X)\leq c_{1}^{\rm CH}(X)\leq c_{1}^{\rm ECH}(X)=c_{Z}(X)$, but we do not know whether equality always holds. This would follow from the stronger conjecture in Remark 1.10.
Acknowledgments.
U.H. acknowledges the support by DFG SFB/TRR 191 ‘Symplectic Structures in Geometry, Algebra and Dynamics’, Projektnummer 281071066-TRR 191. M.H. was partially supported by NSF grant DMS-2005437. V.G.B.R. is grateful for the hospitality of the Institute for Advanced Study, where part of this work was completed. V.G.B.R. was partially supported by NSF grant DMS-1926686, FAPERJ grant JCNE E-26/201.399/2021 and a Serrapilheira Institute grant.